# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v6

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which belongs to the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We can give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5 where we prove that $L_{d}\in\mathcal{NP}$, $A$ is the following:
@@ 8.1 Lower Bounds for Frege Systems
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 8.1 any two Frege systems $p$-simulate each others, we can present the proof of Theorem 7 as follows:
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to produce a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in author’s recent work [Lin21] to show that the language $L_{d}$ is in fact in $\mathcal{NP}$, thus establishing the desired conclusion of this paper.
We stress that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is basically the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since there exists some oracle $A$ such that

### v7

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there exists language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which belongs to the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5 where we prove that $L_{d}\in\mathcal{NP}$, $A$ is the following:
@@ 8.1 Lower Bounds for Frege Systems
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 8.1 we know that any two Frege systems $p$-simulate each others, we then can present the proof of Theorem 7 as follows:
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to produce a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
We stress that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is essentially the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since by the result [BGS75], there exists some oracle $A$ such that